solve by using the quadratic formula
x^2+x-2=0

Answer by uksraj(9)

x^2+x-2=0
the standard form of quadratic equation is
ax^2+bx+c=0
comparing standard form with given equation we get
a=1; b=1 & c=-2
the standard solution is
Solution by SOLVE quadratic equation with variable
Quadratic equation (in our case ) has the following solutons:


For these solutions to exist, the discriminant should not be a negative number.
First, we need to compute the discriminant : .
Discriminant d=9 is greater than zero. That means that there are two solutions: .



Quadratic expression can be factored:
Again, the answer is: 2, -1. Here's your graph:



by factorisation
x^2+2x-x-2=0
x(x+2)-1(x+2)=0
(x-1)(x+2)=0
x=1
x=-2